OpenStudy (anonymous):

Find the inverse of the function. f(x) = x3 - 3

OpenStudy (olivya999):

f−1(x)=x+3−−−−−√3

OpenStudy (anonymous):

Thank you!

RhondaSommer (rhondasommer):

ummm i dont know how you got that but...

RhondaSommer (rhondasommer):

\[x=y^3-3\] -3 -3 \[x-3=y3\] ------ \[\sqrt[3]{x-3}=y\]